OpenStudy (anonymous):

Direction varation The amount of money Sharita makes varies directly with the number of hours she works. If she earns $41.25 for 5 hours of work, how much will she earn if she works 8 hours?

OpenStudy (anonymous):

41.25/5 = x/8 (41.25 for 5 hrs = x/8 hrs) cross multiply (5)(x) = (8)(41.25) 5x = 330 x = 330/5 x = 66

OpenStudy (anonymous):

understand ?

OpenStudy (anonymous):

@kelliegirl33 y 5 first

OpenStudy (anonymous):

@kelliegirl33 can u walk me threw this one Tommy averages one hit for each five times at bat. Use a direct variation formula to determine how many hits would be expected from Tommy after 20 times at bat.

OpenStudy (anonymous):

Tommy averages one hit for each five times at bat. Use a direct variation formula to determine how many hits would be expected from Tommy after 20 times at bat.

OpenStudy (anonymous):

because we are cross multiplying. Take the numerator of one and multiply it by the denominator of the other one and vice versa. Then solve for the missing variable. 1/5 = x/20 (1 hit for 5 times at bat = x hits for 20 times at bat. cross multiply (5)(x) = (20)(1) do you understand so far ?

OpenStudy (anonymous):

can you solve for x now ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

(5)(x) = (20)(1) multiply them 5x = 20 now divide by 5 to get x by itself x = 4 understand ?

OpenStudy (anonymous):

you only set it up this way if it varies directly, not if it varies indirectly

OpenStudy (anonymous):

If it varies indirectly, you have to use inverse proportion